For 2y = - 4 + 24x + 100
when y = a + bx + c
what is the x value of the vertex?

Answer by Hawksfan(61)



The vertex x = -b/2a where these are the coefficients
you must mean 2y = -4x^2 + 24x +100 not just -4
if so
get y by itself by dividing by 2
y = -2x^2 + 12x + 50
now a = -2 and b = 12
x = - (12/2(-2)) = 3 y = -2* 3^2 + 12*3 +50
y = -2 * 9 + 36 + 50 = = 68
Vertex is (6, 68)